OpenStudy (dancergirl45):

\PLEASE WILL FAN AND MEDAL A polynomial has been factored, as shown below: f(x) = (x − 3)(x + 8)(x − 11) What are the zeros of the polynomial? 3, 8, 11 3, −8, 11 −3, 8, −11 −3, 3, −8, 8, −11, 11

OpenStudy (anthonyym):

In other words, the question asks when the polynomial f(x)=0 You have f(x) = (x − 3)(x + 8)(x − 11). Equal that to 0 because that's what the question asks. 0 = (x − 3)(x + 8)(x − 11). Can you solve it now?

OpenStudy (dancergirl45):

so would it be c

OpenStudy (anthonyym):

Also, since the polynomial is multiplied, for f(x) to be 0, either one of the factors has to be 0. So, either x - 3 = 0 x + 8 = 0, or x - 11 = 0

OpenStudy (dancergirl45):

OH ok so then it would be a

OpenStudy (anthonyym):

Yes that's correct

OpenStudy (dancergirl45):

Thank you so much!!

OpenStudy (anthonyym):

No problem!

OpenStudy (phi):

no, it's not choice A

OpenStudy (anthonyym):

oops good catch